# Online Complexity Estimation for Repetitive Scenario Design

### v1

@@ I INTRODUCTION
This paper is concerned with the problem of using the optimal amount of data in repetitive scenario design. Scenario design is a powerful tool for designing “optimal” solutions while guaranteeing that some random constraint $g(x)\leq 0$, where $g:\mathsf{X}\to\mathbb{R}$ is picked randomly, is satisfied with high probability; see, e.g., [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. The principle of scenario design is to replace the random constraint by $N$ i.i.d. samples (called scenarios) of it, namely $g_{i}(x)\leq 0$ for $i=1,\ldots,N$. The final solution is then selected—e.g., by optimizing some preference criterion J(x)J(x)—among all solutions that satisfy the sampled constraints. To fix ideas, let us consider the example of optimal path planning in an uncertain environment consisting of randomly positioned obstacles (see Fig. 1). In this context, the constraint $g(x)\leq 0$ is to avoid the obstacles and its probability distribution is given by the probability distribution of the position of the obstacles (see Fig. 1(a)). The scenario design approach consists in drawing $N$ sampled positions of the obstacles, and finding a path that avoids the obstacles in all—or a predefined fraction—of the sampled positions (see Fig. 1(b)). Under some conditions on the problem, and if $N$ is large enough, probabilistic guarantees on the constraint violation probability (called the risk) of the solution returned by the scenario design algorithm can be given by the theory of scenario design [1, 9, 16, 3, 4, 17].
However, a large sample size $N$ can severely affect the cost of finding a feasible solution to the sampled problem. For instance, if the constraints are nonconvex (as in Fig. 1 for instance), the cost of solving the sampled problem can be exponential in $N$. Moreover, in cases where one can tolerate some probability of failure, sampling too many constraints can lead to overly conservative solutions [4]. Therefore, finding the smallest $N$ that guarantees an upper bound on the risk is of paramount importance for practical applications.
@@
We address this question in the context of repetitive scenario design. Repetitive scenario design is when a similar or slowly-varying scenario design task is performed repeatedly. In other words, it is an online version of scenario design that accounts for continued updates in the data and the environment, assuming small magnitude of the updates. Online path planning is a good example, where path planning (Fig. 1) is performed repeatedly to account for changes in the positions of the obstacles and the initial condition of the system. Because the task is repeated, and under sufficiently slow variations, we can use the information gathered from previous computations to optimize the sample size $N_{t}$ in future steps, so that ultimately the sample size is optimal, i.e., $N_{t}\to N_{\star}$, where $N_{\star}$ is the optimal sample size. Nevertheless, this must be done in a cautious way because we cannot exceed the risk tolerance too often during the task. This precludes naive strategies like “augmenting $N_{t}$ if we exceed the risk tolerance at step t1t-1, and decreasing $N_{t}$ if we meet the risk tolerance at step t1t-1”.
@@ II PROBLEM STATEMENT
To formalize the above, we introduce some notation. Given $\xi\in\Xi$ and $(g_{1},\ldots,g_{N})\in\mathcal{G}^{N}$, we denote by $\mathsf{A}_{\xi}(g_{1},\ldots,g_{N})$ the solution of (3) with $g_{1},\ldots,g_{N}$. [^1] Without loss of generality, we assume that the solution exists and is unique; see, e.g., [3] for ways to handle problems with no solutions or non-unique solutions. Note that this definition of $A$ is fully general: it includes any scenario optimization problems, and any method to obtain an exact or approximate solution to it. Allowing approximate solutions is particularly relevant when dealing with nonconvex or NP-hard optimization problems. Given $x\in\mathsf{X}$ and $\xi\in\Xi$, we denote by $V_{\mathsf{P}}(x,\xi)$ its risk, i.e., the probability of violating the constraint $g(x,\xi)\leq 0$ with respect to $\mathsf{P}$: $V_{\mathsf{P}}(x,\xi)=\mathsf{P}[g(x,\xi)>0]$. When $\mathsf{P}$ and $\xi$ are clear from the context, we write $\mathsf{A}$ and $V(x,\xi)$ for $\mathsf{A}_{\xi}$ and $V_{\mathsf{P}}(x,\xi)$. The probability that the solution of (3) is feasible for (2) is then given by
@@
Let $\mathcal{G}$, $\mathsf{P}$ and $\mathsf{A}$ be as above. Given $\epsilon\in[0,1]$ and $\beta\in[0,1]$, let $N_{\xi}=\min\,\{N\in\mathbb{N}:C_{\xi}(\epsilon,N)\geq\beta\}$. Find a repetitive scenario design algorithm such that, with high probability, $N_{t}\approx N_{\xi_{t}}$ for all $t\in\mathbb{N}_{>0}$ large enough, where $N_{t}$ is the sample size used at step $t$ of the algorithm and $\xi_{t}$ is the value of the external parameter at step $t$.
@@
$\begin{array}[]{cl}\min\limits_{x_{1},\ldots,x_{H}\in\mathsf{B}}&\lVert x_{H}-\xi_{T}\rVert_{2}\\ \text{s.t.}&\lVert x_{t}-x_{t-1}\rVert_{2}\leq\delta,\quad t\in[H],\\ &g(x_{t})\leq 0,\quad t\in[H],\quad g\in\mathcal{G},\end{array}$ (5)
where $x_{0}=\xi_{S}$, $\mathsf{B}=[0,5]\times[0,3]$, and for each $g\in\mathcal{G}$, $g(x)$ is the signed distance (positive in case of collision) between $x$ and the obstacles in one specific position. In the example of Fig. 2(a), each $g\in\mathcal{G}$ has the form
@@
The determination of an optimal sample size $N$ is the central question of this paper.
@@ III-B Distribution shape and fitting
The MLE $\theta_{\star}(\mathcal{D})$ for the problem in Example 1 is represented in Fig. 2(b). We collected $M=1000$ data points by solving the scenario problem 10001000 times with $N=20$ and $H=100$.
@@ V-A Fixed-complexity problems
This problem has fixed complexity $d=20$. The evolution of $N_{t}$ and $\theta_{t}$ over the first $1000$ steps is represented in Fig. 3. We also represented the cumulative of $V(x_{t})$. We notice that $\theta_{t}\to 20$, and $V(x_{t})\leq\epsilon$ with frequency at least $\beta$.
@@ V-B Beyond fixed-complexity problems
We also tried with the probability distribution used in [13, Fig. 9]. The results are given in Fig. 4. In that case, however, the frequency of $V(x_{t})\leq\epsilon$ is smaller than $\beta$. This failure to meet the safety requirements can be explained by the fact that the problem is far from having fixed complexity (see [13, Fig. 9]). In future work, we plan to investigate algorithmic ways to detect such problems that have a non-fixed complexity with high variance, and provide sound ways to converge to their optimal sample size.

### v2

@@ I INTRODUCTION
This paper is concerned with the problem of using the optimal amount of data in repetitive scenario design. Scenario design is a powerful tool for designing “optimal” solutions while guaranteeing that some random constraint $g(x)\leq 0$, where $g:\mathsf{X}\to\mathbb{R}$ is picked randomly, is satisfied with high probability; see, e.g., [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. The principle of scenario design is to replace the random constraint by $N$ i.i.d. samples (called scenarios) of it, namely $g_{i}(x)\leq 0$ for $i=1,\ldots,N$. The final solution is then selected—e.g., by optimizing some preference criterion J(x)J(x)—among all solutions that satisfy the sampled constraints. To fix ideas, let us consider the example of optimal path planning in an uncertain environment consisting of randomly positioned obstacles (see Fig. 1). In this context, the constraint $g(x)\leq 0$ is to avoid the obstacles and its probability distribution is given by the probability distribution of the position of the obstacles (see Fig. 1(a)). The scenario design approach consists in drawing $N$ sampled positions of the obstacles, and finding a path that avoids the obstacles in all—or a predefined fraction—of the sampled positions (see Fig. 1(b)).
Under some conditions on the problem, and if $N$ is large enough, probabilistic guarantees on the constraint violation probability (called the risk) of the solution returned by the scenario design algorithm can be given by the theory of scenario design [1, 9, 16, 3, 4, 17]. However, a large sample size $N$ can severely affect the cost of finding a feasible solution to the sampled problem. For instance, if the constraints are nonconvex (as in Fig. 1 for instance), the cost of solving the sampled problem can be exponential in $N$. Moreover, in cases where one can tolerate some probability of failure, sampling too many constraints can lead to overly conservative solutions [4]. Therefore, finding the smallest $N$ that guarantees an upper bound on the risk is of paramount importance for practical applications.
@@
We address this question in the context of repetitive scenario design. Repetitive scenario design is when a similar or slowly-varying scenario design task is performed repeatedly. In other words, it is an online version of scenario design that accounts for continued updates in the data and the environment, assuming small magnitude of the updates. Online path planning is a good example, where path planning (Fig. 1) is performed repeatedly to account for changes in the position of the obstacles and the initial condition of the system. Because the task is repeated, and under sufficiently slow variations, we can use the information gathered from previous computations to optimize the sample size $N_{t}$ in future steps, so that ultimately the sample size is optimal, i.e., $N_{t}\to N_{\star}(t)$, where $N_{\star}(t)$ is the optimal sample size at time $t$. Nevertheless, this must be done in a cautious way because we cannot exceed the risk tolerance too often during the task. This precludes naive strategies like “augmenting $N_{t}$ if we exceed the risk tolerance at step t1t-1, and decreasing $N_{t}$ if we meet the risk tolerance at step t1t-1”.
@@ II PROBLEM STATEMENT
To formalize the above, we introduce some notation. Given $\xi\in\Xi$ and $(g_{1},\ldots,g_{N})\in\mathcal{G}^{N}$, we denote by $\mathsf{A}_{\xi}(g_{1},\ldots,g_{N})$ the solution of (3) with $g_{1},\ldots,g_{N}$. [^1] Without loss of generality, we assume that the solution exists and is unique; see, e.g., [3] for ways to handle problems with no solutions or non-unique solutions. Note that this definition of $A$ is very general: it includes any scenario optimization problems, and any method to obtain an exact or approximate solution to it. Allowing approximate solutions is particularly relevant when dealing with nonconvex or NP-hard problems. Given $x\in\mathsf{X}$ and $\xi\in\Xi$, we denote by $V_{\mathsf{P}}(x,\xi)$ its risk, i.e., the probability of violating the constraint $g(x,\xi)\leq 0$ with respect to $\mathsf{P}$: $V_{\mathsf{P}}(x,\xi)=\mathsf{P}[g(x,\xi)>0]$. When $\mathsf{P}$ and $\xi$ are clear from the context, we write $\mathsf{A}$ and $V(x)$ for $\mathsf{A}_{\xi}$ and $V_{\mathsf{P}}(x,\xi)$. The probability that the solution of (3) is feasible for (2) is then given by
@@
Let $\mathcal{G}$, $\mathsf{P}$ and $\mathsf{A}$ be as above. Given $\epsilon\in[0,1]$ and $\beta\in[0,1]$, let $N_{\star}(\xi)=\min\,\{N\in\mathbb{N}:C_{\xi}(\epsilon,N)\geq\beta\}$. Find a repetitive scenario design algorithm such that, with high probability, $N_{t}\approx N_{\star}(\xi_{t})$ for all $t\in\mathbb{N}_{>0}$ large enough, where $N_{t}$ is the sample size used at step $t$ of the algorithm and $\xi_{t}$ is the value of the external parameter at step $t$.
@@
$\begin{array}[]{cl}\min\limits_{x_{1},\ldots,x_{H}\in\mathsf{B}}&\lVert x_{H}-\xi_{T}\rVert_{2}\\ \text{s.t.}&\lVert x_{t}-x_{t-1}\rVert_{2}\leq\delta,\quad t\in[H],\\ &g(x_{t},\xi)\leq 0,\quad t\in[H],\quad g\in\mathcal{G},\end{array}$ (5)
where $x_{0}=\xi_{S}$, $\mathsf{B}=[0,5]\times[0,3]$, and for each $g\in\mathcal{G}$, $g(x,\xi)$ is the signed distance (positive in case of collision) between $x$ and the obstacles in one specific position. In the example of Fig. 2(a), each $g\in\mathcal{G}$ has the form
@@
The determination of an optimal sample size $N$ is the central question of this paper. $\triangleleft$
@@ III-B Distribution shape and fitting
The MLE $\theta_{\star}(\mathcal{D})$ for the problem in Example 1 is represented in Fig. 2(b). We collected $M=1000$ data points by solving the scenario problem 10001000 times with $N=20$ and $H=100$. $\triangleleft$
@@ V-A Fixed-complexity problems
This problem has fixed complexity $d=20$. The evolution of $N_{t}$ and $\theta_{t}$ over the first $1000$ steps is represented in Fig. 3. We also represented the cumulative of $V(x_{t})$. We notice that $\theta_{t}\to 20$, and $V(x_{t})\leq\epsilon$ with frequency almost $\beta$.
@@ V-B Beyond fixed-complexity problems
We also tried with the probability distribution used in [13, Fig. 9]. The results are given in Fig. 4. In this case, however, the frequency of $V(x_{t})\leq\epsilon$ is smaller than $\beta$. This failure to meet the safety requirements can be explained by the fact that the problem is far from having fixed complexity (see [13, Fig. 9]). In future work, we plan to investigate algorithmic ways to detect such problems that have a non-fixed complexity with high variance, and provide sound ways to converge to their optimal sample size.